a refrigerating plant is required to produce 2.5 tonnes of ice per day at -4°c from water at 20°c if the temperature range in the compressor is between 25°c and -6°c calculate power required to drive the compressor latent heat of ice= 335 kj/kg and specific heat of ice = 2.1 kj/kg k


1
Expert's answer
2021-06-22T05:10:21-0400

"m = 896.55 \frac {kg}{sec}"


"h_2 = 335 \frac{kj} {kg}"


"h_1 = 264.68 \frac {kJ}{kg}" ...From refrigeration Chart


Power required to drive compressor "= m(h_2-h_1) = 896.55(335-264.68) = 63kW"